Problem A.2

Let f(x), g(x) and h(x) be the functions from Problem A.1. Find the derivative λ′(x) of the following function with respect to x:

λ(x) = f(x) · g(x) + f(x) · h(x) − g(x) · h(x)


1
Expert's answer
2021-11-22T16:35:26-0500

"\lambda(x) = f(x) · g(x) + f(x) · h(x) − g(x) · h(x) \\
Using \ the \ product \ rule \ to \ get \ derivative \ \lambda'(x) \\
\therefore \lambda'(x) = \\ 
f(x) · g'(x) + f'(x) · g(x) \\ 
+ f(x) · h'(x) + f'(x) · h(x) \\ 
− g(x) · h'(x) + g'(x) · h(x) \\

From \ Problem\ A.1, \\
f(x) = x - 2 \\
g(x) = 4 - (x - 6)^2 + 2 \\ 
h(x) = x - 6 \\
\therefore \lambda(x) = \\ 
(x - 2) · (4 - (x - 6)^2 + 2) + (x - 2) · (x - 6) \\ 
- (4 - (x - 6)^2 + 2) · (x - 6) \\

f'(x) = \frac{d} {dx} (x-2) = 1 \\
g'(x) = \frac{d} {dx} (4-(x-6)^2+2) = -2(x-6) \\
h'(x) = \frac{d} {dx} (x-6) = 1\\

\lambda'(x) = \\ 
[(x-2) · (-2(x-6)) + 1 · (4-(x-6)^2 + 2)] \\
+ [(x-2) · (1) + 1 · (x-6)] \\
- [(4-(x-6)^2 +2) · (1) + (-2 (x-6)) · (x-6)] \\

\lambda'(x) = \\
(x-2)(-2x+12) + (4-(x-6)^2 +2) \\
+ (x-2) + (x-6) \\ 
- (4 - x^2 + 12x - 36 + 2 - 2x^2 + 24x - 72) \\

\lambda'(x) = \\
-2x^2 + 12x + 4x - 24 + 4 - x^2 + 12x - 36 \\ 
+ 2 + x - 2 + x - 6 \\
- 4 + x^2 - 12x + 36 - 2 + 2x^2 - 24x + 72 \\

\lambda'(x) = -6x + 40 \\
\lambda'(x) = -2(3x - 20)"